# How Combined Pairwise and Higher-Order Interactions Shape Transient Dynamics

### v1

@@ II.3.3 Analytical Insights through Linear Theory
To further validate our results analytically, we employ linear algebraic theories. We use local stability analysis and examine the leading eigenvalues of the Jacobian matrix. The leading eigenvalue is defined as the eigenvalue with the largest negative real part. The following table compares the leading eigenvalues for the system using the same parameter set as in Fig. 2.
@@
In a small $\epsilon$-neighborhood with $\epsilon>0$ near the equilibrium point, the solution to linearised Eq. 1 for $N=3$ can be written as:
@@
Since $\mathscr{R}$ is negative in our non-hyberbolic system, the term $\exp(\mathscr{R}t)$ represents an exponential decay. Although one of the eigenvalues is zero in our case, indicating the corresponding term $C_{1}v_{1}$ neither decays nor grows, which prevents a conclusive determination of linear stability, we establish global stability through the construction of a suitable Lyapunov function. We can assert that the magnitude of $\mathscr{R}$ governs the rate of convergence towards equilibrium, with the frequency of oscillation determined by the imaginary part $\Im$. However, similar to how linear stability analysis is restricted to the vicinity of the equilibrium, the estimation of the convergence rate based on the real part of the leading eigenvalue $\mathscr{R}$ is only valid within a local neighborhood. Nonetheless, our analysis, grounded in the real part of the leading eigenvalues, offers a robust method for estimating the rate at which populations converge to coexistence abundances. Irrespective of the initial abundances, i.e., even if the populations begin at vastly different levels, the system will reliably move toward a balanced state of coexistence, where each species maintains a stable abundance, and our theoretical predictions regarding the rate of convergence towards the coexistence equilibrium align with observed dynamics in the system. This behavior is validated by the results shown in Table 1 and Fig. 4.
@@
Figure 5: In this analysis, we explore how different interaction structures affect the transient times at varying levels of higher-order interactions. The boxplot shows the variation in transient times for different values of $\gamma_{2}$ across $20$ randomly generated interaction matrices $A$, with a fixed initial condition $(0.15,0.35,0.5)$. The orange dots represent mean transient times, while the yellow line marks the median. As $\gamma_{2}$ increases from $0.05$ (minimal higher-order interactions) to $1.0$ (solely higher-order interactions), transient times decrease, demonstrating the structural robustness of the system in reaching equilibrium. Notably, the coexistence equilibrium remains unchanged across varying $A$, emphasizing that the choice of the particular circulant matrix preserves the coexistence equilibrium and higher-order interactions accelerate convergence to the ecological stability.
@@ II.3.4 Effects of Varying Interaction Matrices with Fixed Initial Species Abundances
Up to this point, we have explored the influence of higher-order interactions by varying the initial conditions while keeping the interaction matrix $A$ fixed. Alternatively, one can take the opposite approach by keeping the initial condition constant and instead varying the interaction matrix $A$. To explore how the structure of interaction impacts transient dynamics, we fix an initial condition at $(0.15,0.35,0.5)$, ensuring $x_{1}(0)+x_{2}(0)+x_{3}(0)=1.0$, and vary the interaction matrix $A$ by randomly sampling 20 values of $\alpha$ from a uniform distribution over $(0,1)\setminus{0.5}$. For each $\alpha$, we generate $20$ different circulant $A$ matrices for $N=3$ and compute the mean transient time, which is plotted in Fig. 5. We observe that increasing $\gamma_{2}$ consistently reduces transient time, confirming the influence of higher-order interactions on faster convergence. Notably, while varying the initial conditions slightly impacts the transient dynamics, the structure of interactions plays a far more dominant role in determining the transient times, as indicated by the significantly higher variance in transient times compared to different initial conditions. This suggests that the ecological architecture, particularly how species interact through both pairwise and higher-order interactions, exerts a stronger influence on system behavior than initial population distributions. Despite the changes in interaction matrices $A$, the coexistence equilibrium remains constant, underscoring the resilience of the equilibrium point to structural variations in the interaction network. Our circulant matrices ensure that ecosystems with varying circulant interaction patterns will still converge to the same stable coexistence equilibrium. This highlights the role of symmetry and uniformity in maintaining stable ecological balances across diverse interaction networks. However, it should be noted that higher-order interactions accelerate convergence to equilibrium regardless of the interaction matrix structure Grilli et al. (2017).
Figure 6: (a) The subfigure illustrates how transient time varies with different values of $\gamma_{2}\in[0,1]$ and $\gamma_{3}\in[0,1]$ with $\gamma_{1}+\gamma_{2}+\gamma_{3}=1$, representing the influence of first and second higher-order interactions (HOIs) on system dynamics. The y-axis shows a more pronounced reduction in transient time along the $\gamma_{3}$ axis compared to the $\gamma_{2}$ axis, indicating that second HOI have a stronger impact on accelerating convergence to the equilibrium. This result reflects the ecological insight that more complex interactions (higher-order) promote faster stabilization of ecosystems, enhancing their ability to quickly reach a stable coexistence state. The asymmetry observed between $\gamma_{2}$ and $\gamma_{3}$ underscores the differential effects of various HOIs on the system’s transient dynamics. (b) The subfigure illustrates how the real part of the leading eigenvalue shifts with varying values of $\gamma_{2}\in[0,1]$ and $\gamma_{3}\in[0,1]$, constrained by $\gamma_{1}+\gamma_{2}+\gamma_{3}=1$, capturing the convergence rate near the system’s equilibrium. The $\gamma_{3}$ axis reveals a more significant negative real part of the eigenvalue compared to the $\gamma_{2}$ axis, indicating that second HOI have a stronger influence on accelerating the system’s return to equilibrium. This asymmetry between $\gamma_{2}$ and $\gamma_{3}$ reflects the different biological roles these interactions might play in stabilizing species coexistence; for instance, second HOI could more effectively counterbalance competitive imbalances in ecological communities. This result is consistent with subfigure (a), reinforcing the distinct contributions of first- and second HOI to the system’s transient dynamics.
@@ II.3.5 Exploring Beyond First Higher Order Interactions
Now we analytically compute the Jacobian matrix for $N=5$, incorporating both first $(\gamma_{2})$ and second $(\gamma_{3})$ higher-order interactions (HOIs) alongside pairwise interactions $(\gamma_{1})$. Although the eigenvalues derived from this analysis are complicated and difficult to express explicitly, we gain better insight into their behavior by plotting them in a two-dimensional triangular space defined by the constraints $\gamma_{2}=0$, $\gamma_{3}=0$, and $\gamma_{2}+\gamma_{3}=1$. This analysis aligns with our theoretical expectations; despite the non-hyperbolic nature of the model, the real part of the leading eigenvalue effectively reveals the system’s convergence rate toward the coexistence equilibrium. As shown in Fig. 6(b), there is a clear asymmetry, with $\gamma_{3}$ exerting a stronger influence on the convergence dynamics than $\gamma_{2}$. It is important to note that while the transient time decays more rapidly than exponentially, the real part of the leading eigenvalue exhibits a linear trend. Though both behaviors reflect the rate of convergence toward equilibrium, there is a subtle distinction between the patterns observed in the two subfigures. This behavior is consistent with the earlier observation in Fig. 4.
@@ III Conclusion & Outlook
Our findings reveal that increasing the proportion of higher-order interactions accelerates the convergence to this equilibrium. This is due to higher-order interactions promoting a more efficient stabilization process, whereas pairwise interactions tend to prolong the transient phase, thereby slowing convergence. Although the linearized system with a combination of pairwise and higher-order interactions possesses a zero eigenvalue, the negative real parts of the leading eigenvalues indicate that higher-order interactions contribute to faster convergence. This suggests that, biologically, systems with a greater fraction of higher-order interactions are more adept at reaching stable states, whereas pairwise interactions alone may impede this process by increasing transient dynamics.
Our framework offers fresh insights into natural systems where extended transient phases are prevalent. By integrating both pairwise and higher-order interactions, our model not only deepens the understanding of dynamic stabilization but also suggests new mechanisms for addressing prolonged transient behaviors. These concepts could pave the way for identifying and developing strategies to manage and mitigate such dynamics in ecological and evolutionary systems. A potential avenue for future research is to explore how transient dynamics are affected by time-varying interactions Nag Chowdhury et al. (2019b); Nag Chowdhury and Ghosh (2019); Nag Chowdhury et al. (2020a, b), a phenomenon commonly observed in natural systems. This remains a central and intriguing direction for future research, with the potential to uncover a broad range of fascinating outcomes.
@@ acknowledgement
SC would like to acknowledge Prof. Soumitra Banerjee for his deep insights, invaluable comments on the work during his IISER Kolkata days. SC would also Dr. Chittaranjan Hens for his valuable discussions during his lab visit at IIIT Hyderabad funded by his grant at IIIT and for his strong encouragement in pursuing this research. SNC gratefully acknowledges the support of the National Science Foundation under Grant No 1840221 during his postdoctoral tenure at UC Davis, where the problem was first discussed with SC. Further in-depth discussions with SC took place during the 7th International Conference on Complex Dynamical Systems and Applications 2024, held on January 25-27, 2024, at the Digha Science Centre & National Science Camp, India, while SNC was a visiting scientist at the Physics and Applied Mathematics Unit of the Indian Statistical Institute, Kolkata. SNC expresses gratitude to Prof. Ulrike Feudel for the opportunity to discuss this topic during this period. SNC would also like to extend his gratitude to Constructor University for its support during his postdoctoral tenure, during which the final stages of this manuscript were completed. SNC would also like to express his gratitude to Prof. Syamal K. Dana for his valuable discussions during the Dynamics Days Europe 2024, held from July 29 to August 2, 2024, at Constructor University Bremen gGmbH, Germany, and for his encouragement in pursuing this research. SNC would like to thank Prof. Alan Hastings for his insightful discussions on this topic during his stay at UC Davis. Special thanks also go to his wife, Dr. Srilena Kundu, for her valuable suggestions, which significantly improved the manuscript.

### v2

@@ II.3.3 Analytical Insights through Linear Theory
To further validate our results analytically, we employ linear algebraic theories. We use local stability analysis and examine the leading eigenvalues of the Jacobian matrix associated with the linearized system at the coexistence equilibrium. The leading eigenvalue is defined as the eigenvalue with the largest negative real part. The following table compares the leading eigenvalues for the system using the same parameter set as in Fig. 2.
@@
In a small $\epsilon$-neighborhood with $\epsilon>0$ near the equilibrium point, the solution to Eq. 1, i.e, the solution to linearised Eq. 1 for $N=3$ can be written as:
@@
Since $\mathscr{R}$ is negative in our non-hyberbolic system, the term $\exp(\mathscr{R}t)$ represents an exponential decay. Although one of the eigenvalues is zero in our case, indicating the corresponding term $C_{1}v_{1}$ neither decays nor grows, which prevents a conclusive determination of linear stability, we establish global stability through the construction of a suitable Lyapunov function. We can assert that the magnitude of $\mathscr{R}$ governs the rate of convergence towards equilibrium, with the frequency of oscillation determined by the imaginary part $\Im$. However, similar to how linear stability analysis is restricted to the vicinity of the equilibrium, the estimation of the convergence rate based on the real part of the leading eigenvalue $\mathscr{R}$ is only valid within a local neighborhood of the equilibrium. Nonetheless, our analysis, grounded in the real part of the leading eigenvalues, offers a robust method for estimating the rate at which populations converge to coexistence abundances. Irrespective of the initial abundances, i.e., even if the populations begin at vastly different levels, the system will reliably move toward a balanced state of coexistence, where each species maintains a stable abundance, and our theoretical predictions regarding the rate of convergence towards the coexistence equilibrium align with observed dynamics in the system. This behavior is validated by the results shown in Table 1 and Fig. 4.
@@
Figure 5: The influence of higher-order interactions ($\gamma_{2}$) on transient times across different interaction matrices ($A$). In this analysis, we explore how different interaction structures affect the transient times at varying levels of higher-order interactions. The boxplot shows the variation in transient times for different values of $\gamma_{2}$ across $20$ randomly generated interaction matrices $A$, with a fixed initial condition $(0.15,0.35,0.5)$. The red dots represent mean transient times, while the yellow line marks the median. As $\gamma_{2}$ increases from $0.05$ (minimal higher-order interactions) to $1.0$ (solely higher-order interactions), transient times decrease, demonstrating the structural robustness of the system in reaching equilibrium. Notably, the coexistence equilibrium remains unchanged across varying $A$, emphasizing that the choice of the particular circulant matrix preserves the coexistence equilibrium and higher-order interactions accelerate convergence to the ecological stability.
Figure 6: The impact of first-order ($\gamma_{2}$) and second-order ($\gamma_{3}$) higher-order interactions (HOIs) on transient times. (a) The subfigure illustrates how transient time varies with different values of $\gamma_{2}\in[0,1]$ and $\gamma_{3}\in[0,1]$ with $\gamma_{1}+\gamma_{2}+\gamma_{3}=1$, representing the influence of first and second higher-order interactions (HOIs) on system dynamics. The y-axis shows a more pronounced reduction in transient time along the $\gamma_{3}$ axis compared to the $\gamma_{2}$ axis, indicating that second HOI have a stronger impact on accelerating convergence to the equilibrium. This result reflects the ecological insight that more complex interactions (higher-order) promote faster stabilization of ecosystems, enhancing their ability to quickly reach a stable coexistence state. The asymmetry observed between $\gamma_{2}$ and $\gamma_{3}$ underscores the differential effects of various HOIs on the system’s transient dynamics. (b) The subfigure illustrates how the real part of the leading eigenvalue of the Jacobian matrix associated with the linearized system at the coexistence equilibrium shifts with varying values of $\gamma_{2}\in[0,1]$ and $\gamma_{3}\in[0,1]$, constrained by $\gamma_{1}+\gamma_{2}+\gamma_{3}=1$, capturing the convergence rate near the system’s equilibrium. The $\gamma_{3}$ axis reveals a more significant negative real part of the leading eigenvalue compared to the $\gamma_{2}$ axis, indicating that second HOI have a stronger influence on accelerating the system’s return to equilibrium. This asymmetry between $\gamma_{2}$ and $\gamma_{3}$ reflects the different biological roles these interactions might play in stabilizing species coexistence; for instance, second HOI could more effectively counterbalance competitive imbalances in ecological communities. This result is consistent with subfigure (a), reinforcing the distinct contributions of first- and second HOI to the system’s transient dynamics.
@@ II.3.4 Effects of Varying Interaction Matrices with Fixed Initial Species Abundances
Up to this point, we have explored the influence of higher-order interactions by varying the initial conditions while keeping the interaction matrix $A$ fixed. Alternatively, one can take the opposite approach by keeping the initial condition constant and instead varying the interaction matrix $A$. To explore how the structure of interaction impacts transient dynamics, we fix an initial condition randomly at $(0.15,0.35,0.5)$, ensuring $x_{1}(0)+x_{2}(0)+x_{3}(0)=1.0$, and vary the interaction matrix $A$ by randomly sampling 20 values of $\alpha$ from a uniform distribution over $(0,1)\setminus{0.5}$. For each $\alpha$, we generate $20$ different circulant $A$ matrices for $N=3$ and compute the mean transient time, which is plotted in Fig. 5. We observe that increasing $\gamma_{2}$ consistently reduces transient time, confirming the influence of higher-order interactions on faster convergence. Notably, while varying the initial conditions slightly impacts the transient dynamics, the structure of interactions plays a far more dominant role in determining the transient times, as indicated by the significantly higher variance in transient times compared to different initial conditions. This suggests that the ecological architecture, particularly how species interact through both pairwise and higher-order interactions, exerts a stronger influence on system behavior than initial population distributions. Despite the changes in interaction matrices $A$, the coexistence equilibrium remains constant, underscoring the resilience of the equilibrium point to structural variations in the interaction network. Our circulant matrices ensure that ecosystems with varying circulant interaction patterns will still converge to the same stable coexistence equilibrium. This highlights the role of symmetry and uniformity in maintaining stable ecological balances across diverse interaction networks. However, it should be noted that higher-order interactions accelerate convergence to equilibrium regardless of the interaction matrix structure Grilli et al. (2017).
@@ II.3.5 Exploring Beyond First Higher Order Interactions
Now we analytically compute the Jacobian matrix for $N=5$, incorporating both first $(\gamma_{2})$ and second $(\gamma_{3})$ higher-order interactions (HOIs) alongside pairwise interactions $(\gamma_{1})$. Although the eigenvalues derived from this analysis are complicated and difficult to express explicitly, we gain better insight into their behavior by plotting them in a two-dimensional triangular space defined by the constraints $\gamma_{2}=0$, $\gamma_{3}=0$, and $\gamma_{2}+\gamma_{3}=1$. This analysis aligns with our theoretical expectations; despite the non-hyperbolic nature of the model, the real part of the leading eigenvalue effectively reveals the system’s convergence rate toward the coexistence equilibrium. As shown in Fig. 6(b), there is a clear asymmetry, with $\gamma_{3}$ exerting a stronger influence on the convergence dynamics than $\gamma_{2}$. It is important to note that while the transient time decays more rapidly in a non-linear fashion maintaining an inverse relationship with $\gamma_{2}$, the real part of the leading eigenvalue exhibits a linear trend for a fixed $\gamma_{3}$. Though both behaviors reflect the rate of convergence toward equilibrium, there is a subtle distinction between the patterns observed in the two subfigures. This behavior is consistent with the earlier observation in Fig. 4.
@@ III Conclusion & Outlook
Our findings reveal that increasing the proportion of higher-order interactions accelerates the convergence to this equilibrium. This is due to higher-order interactions promoting a more efficient stabilization process, whereas pairwise interactions tend to prolong the transient phase, thereby slowing convergence. Although the linearized system with a combination of pairwise and higher-order interactions possesses a zero eigenvalue, the negative real parts of the leading eigenvalues of the Jacobian matrix associated with the linearized system at the coexistence equilibrium indicate that higher-order interactions contribute to faster convergence. This suggests that, biologically, systems with a greater fraction of higher-order interactions are more adept at reaching stable states, whereas pairwise interactions alone may impede this process by increasing transient dynamics.
This study employs a simplified replicator model to explore species interactions and their impact on biodiversity, acknowledging that while it does not capture all complexities of real-world systems, it remains valuable for examining various scenarios. Despite extensive research on replicator dynamics, our understanding of transient dynamics is limited due to a lack of established mathematical techniques. We aim to shed light on ecosystem behavior during transient periods, providing insights that may not only enhance our knowledge of ecosystem dynamics but also suggest pathways for more effective management and conservation practices. We focus on a simple replicator model, employing convex combinations to preserve the coexistence equilibrium while comparing transient times. Our model is biologically relevant, as it operates within the interval $[0,1]$ and reliably converges to the coexistence equilibrium. For simulations, we exclusively use circulant matrices that do not affect the coexistence equilibrium, allowing for meaningful comparisons of transient times. Driven by ecological considerations, we selected this model to explore how different interaction types influence transient dynamics. Our findings highlight the critical role of higher-order interactions in promoting quicker convergence to stable coexistence, which is essential for understanding adaptation and resistance in ecological systems.
Our framework offers fresh insights into natural systems where extended transient phases are prevalent. By integrating both pairwise and higher-order interactions, our model not only deepens the understanding of dynamic stabilization but also suggests new mechanisms for addressing prolonged transient behaviors. These concepts could pave the way for identifying and developing strategies to manage and mitigate such dynamics in ecological and evolutionary systems. We would like to clarify that our findings may not generalize to chaotic or more complex systems, as our analysis is grounded in linear and global stability methods, which are not suited for capturing chaotic dynamics. Likewise, the use of Lyapunov functions is limited to assessing equilibrium stability. We acknowledge that systems exhibiting chaotic behavior require alternative mathematical approaches, which are beyond the scope of this study. Exploring these complexities could offer valuable insights and represents a promising avenue for future research. Another potential avenue for future research is to explore how transient dynamics are affected by time-varying interactions Nag Chowdhury et al. (2019b); Ansarinasab et al. (2024a); Nag Chowdhury and Ghosh (2019); Mikaberidze et al. (2024); Dayani et al. (2023); Nag Chowdhury et al. (2020b, c); Ansarinasab et al. (2024b), a phenomenon commonly observed in natural systems. This remains a central and intriguing direction for future research, with the potential to uncover a broad range of fascinating outcomes.
@@ acknowledgement
SC gratefully acknowledges Prof. Soumitra Banerjee for his deep insights and invaluable feedback during his time at IISER Kolkata, as well as Dr. Chittaranjan Hens for valuable discussions during SC’s lab visit to IIIT Hyderabad, funded by IIIT, and for his strong encouragement in pursuing this research. SNC thanks the National Science Foundation (Grant No. 1840221) for its support during his postdoctoral tenure at UC Davis, where the problem was first discussed with SC. Further in-depth discussions occurred during the 7th International Conference on Complex Dynamical Systems and Applications 2024 in Digha, India, while SNC was a visiting scientist at the Physics and Applied Mathematics Unit, Indian Statistical Institute, Kolkata. Gratitude is extended to Prof. Ulrike Feudel for the opportunity to discuss this work during this period. SNC also thanks Constructor University for its support during his postdoctoral tenure, during which the manuscript’s final stages were completed, and Prof. Syamal K. Dana for his valuable discussions at Dynamics Days Europe 2024 in Bremen, Germany, and his encouragement. Additionally, SNC expresses deep thanks to Prof. Alan Hastings for his insightful discussions at UC Davis and to his wife, Dr. Srilena Kundu, for her valuable suggestions, which significantly improved the manuscript.